A lamp in a childs Halloween costume flashes based on an RC discharge of a capacitor through its resistance. The effective duration of the flash is 0.120 s, during which it produces an average 0.460 W from an average 3.00 v (a) What energy does it dissipate? (b) How much charge moves through the lamp? (c) Find the capacitance. (d) What is the resistance of the lamp?
Answer #1

Solution,

A)

Energy = power x time = 0.460 x 0.12 = 0.055 J

B)

Power, P = voltage x current

0.460 = 3 x I

Current, I = 0.153 A

Charge, Q = I t = 0.153 x 0.12 = 0.018 C

C)

Charge, Q = CV

0.018 = C x 3

C = 0.0061 F

D)

Resistance, R = V/I = 3/0.153 = 19.61 ohm

Comment in case any doubt please rate my answer ...